OpenStudy (anonymous):

Given that f(s)=4s^2+3s+8 Find the ratio (f(t+h)-f(t))/h

OpenStudy (johnweldon1993):

Should that original equation be f(t) = 4s² + 3s + 8 ? *just making sure it is f(t)*

OpenStudy (anonymous):

no its f(s)

OpenStudy (johnweldon1993):

Well I'm going to go on assuming that it is... you have f(t) = 4t² + 3t + 8 And you are finding the derivative using the definition ... \[\frac{ f(t + h) - f(t) }{ h }\] What you do...is take your equation....and start off like this \[\frac{ 4(t + h)^2 + 3(t + h) + 8 - f(t) }{ h }\] Notice that I replaced all the 'x' in the original equation with the new (x + h) Now you finish it off with that - f(t) all you do now is write in the original equation again since that IS f(t)..so \[\frac{ 4(t + h)^2 + 3(t + h) + 8 - (4t² + 3t + 8) }{ h }\] Now just simplify through...

OpenStudy (johnweldon1993):

So we'll have \[\frac{ 4t^2 + 8th + 4h^2 + 3t + 3h + 8 - 4t^2 - 3t - 8 }{ h }\] And now we just combine like terms to simplify \[\frac{4h^2 + 8th + 3h }{ h }\] Now we can divide through by the 'h' to get \[4h + 8t + 3\] We know that 'h' is approaching 0 so f'(t) = 8t + 3

OpenStudy (anonymous):

Thank u :)